# Revisiting, Benchmarking and Understanding Unsupervised Graph Domain Adaptation

### v1

@@ Abstract
Unsupervised Graph Domain Adaptation (UGDA) involves the transfer of knowledge from a label-rich source graph to an unlabeled target graph under domain discrepancies. Despite the proliferation of methods designed for this emerging task, the lack of standard experimental settings and fair performance comparisons makes it challenging to understand which and when models perform well across different scenarios. To fill this gap, we present the first comprehensive benchmark for unsupervised graph domain adaptation named GDABench, which encompasses 16 algorithms across 5 datasets with 74 adaptation tasks. Through extensive experiments, we observe that the performance of current UGDA models varies significantly across different datasets and adaptation scenarios. Specifically, we recognize that when the source and target graphs face significant distribution shifts, it is imperative to formulate strategies to effectively address and mitigate graph structural shifts. We also find that with appropriate neighbourhood aggregation mechanisms, simple GNN variants can even surpass state-of-the-art UGDA baselines. To facilitate reproducibility, we have developed an easy-to-use library PyGDA for training and evaluating existing UGDA methods, providing a standardized platform in this community. Our source codes and datasets can be found at https://github.com/pygda-team/pygda.
@@ 1 Introduction
The last decade has witnessed significant advancements in Graph Neural Networks (GNNs) with their successful applications spanning various fields [1, 2], including social network analysis [3], protein interaction prediction [4], and traffic flow forecasting [5], etc. However, the presence of distribution shifts [6] and label scarcity in real-world graph data impedes the ability of existing GNN models to adapt to new domains [7]. To addresses this challenge, Unsupervised Graph Domain Adaptation (UGDA) has become an important solution for transferring knowledge from a labeled source graph to an unlabeled target graph. This powerful paradigm has been widely studied, unlocking broader application for graph neural networks.
@@
To fill these gaps, we revisit existing UGDA algorithms and conduct a comprehensive benchmark named GDABench. Specifically, GDABench includes 16 state-of-the-art UGDA models and 5 diverse graph datasets covering node attributes, graph structures, and label proportion shifts, resulting in 74 different adaptation tasks. Additionally, we also explore the limits of GNN transferability by combining 7 GNN variants with 2 domain alignment and 3 unsupervised graph learning techniques. Our work is the first to provide a rigorous empirical analysis of how various aggregation mechanisms influence alignments in domain adaptation task. Through comprehensive experiments, we observe that: (1) the performance of current UGDA models varies greatly across different datasets and adaptation scenarios; (2) it is crucial to develop tailored strategies to address graph structural shifts, especially when the distribution discrepancies are significant; (3) the GNN’s transferability in UGDA heavily relies on two factors: aggregation scope and aggregation architecture, which are influenced by the severity of label shift and the level of graph heterophily, etc; (4) the inherent adaptability of GNNs is largely underestimated by existing methods, which motivates the investigate of a simple yet effective model that relies heavily on GNN’s property. More insights can be found in Section 5.
@@
  • We introduce GDABench, the first comprehensive benchmark for unsupervised graph domain adaptation. It includes 16 recent state-of-the-art methods across 5 real-world datasets with diverse range of adaptation tasks.
@@ 2.1 Problem Definition
Given an attributed graph $G=(\mathcal{V},\mathcal{E})$ with $n$ nodes and $m$ edges, the node attribute matrix is represented as $\mathbf{X}=\{{x_{v}|v\in}\mathcal{V}\}\in\mathbb{R}^{n\times d}$, where $d$ is the dimension of node attributes. The adjacency matrix is denoted as $\mathbf{A}\in\mathbb{R}^{n\times n}$, where $\mathbf{A}_{i,j}=1$ means there exists an edge $e_{i,j}\in\mathcal{E}$ connecting node $v_{i}$ and $v_{j}$, and $\mathbf{A}_{i,j}=0$ otherwise. $\mathbf{Y}\in\mathbb{R}^{N\times C}$ denotes the node label matrix, where $C$ is the category of node labels. We focus on Unsupervised Graph Domain Adaptation (UGDA) on node classification tasks. It is a non-trivial task due to the complex domain shift between source and target graphs. Formally, given a labeled source graph $\mathcal{G}_{S}=(\mathcal{V}_{S},\mathcal{E}_{S},\mathcal{Y}_{S})$ and an unlabeled target graph $\mathcal{G}_{T}=(\mathcal{V}_{T},\mathcal{E}_{T})$ with the data shift that $\mathbb{P}_{S}(G)\neq\mathbb{P}_{T}(G)$. The goal is to train a graph neural network model $h:\mathcal{G}\rightarrow\mathcal{Y}$ using the labeled source graph $\mathcal{G}^{s}$ and the unlabeled target graph $\mathcal{G}^{t}$ to accurately predict labels on target graph $\mathcal{G}^{t}$.
@@ 2.2 Related Work
Classical Methods. There are two classical ways to minimize the domain discrepancy [32, 33]: minimizing pre-defined probability discrepancy metrics [34, 35] and adversarial learning techniques [36, 37, 38, 39]. For the pre-defined probability discrepancy metrics minimization models, node representations are first obtained from the encoder, and then domain-invariant representations are learned by minimizing probability discrepancy distances, such as MMD [40], CMD [35], etc. Instead of explicitly minimizing domain discrepancies, some methods incorporate the encoder with a domain classifier that predicts the domain from which the representation originates. For example, DANN [36] facilitates domain invariant learning (DIRL) to distinguish between source and target samples in the latent space. Building upon this framework, WDGRL [41] replaces the domain classifier with a network that learns an approximate Wasserstein distance. These classical methods are designed for computer vision and natural language processing tasks, where samples are independently and identically distributed [42]. However, marginal alignment of node representations in non-graph DA research is insufficient for graph-structured data, of which the distribution shift becomes more complex due to the interconnection among different nodes.
@@
Specialized Methods for Graph Data. To tackle the unique challenges of graph knowledge transfer, several approaches have been proposed [8, 10, 43, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
@@
The dominant idea of existing work on node-level UGDA focused on leveraging node representations as intermediaries and then minimizing the domain shift by adversarial learning. This kind of methods incorporate classical DA methods with deep node embedding by designing a special encoder to learn transferable features. DANE [9] uses shared weight GCNs to get node representations and then handles distribution shift via least square generative adversarial network. ACDNE [10] utilizes two feature extractors to jointly preserve attributed affinity and topological proximities as deep network embedding module and incorporates a domain classifier to make node representations label-discriminative. UDAGCN [11] develops a dual graph convolutional network with attention mechanism to jointly exploit local and global consistency for effective graph representation learning. ASN [43] separates domain-private and domain-shared information and combines local and global consistency to capture network topology information. AdaGCN [12] leverages GCN to integrate network topology and combines adversarial domain adaptation with graph convolution. CWGCN [21] puts forward a two-step correntropy-induced Wasserstein GCN, which first suppresses the noisy nodes in the source graph and then learns the target GCN based on extending the Wasserstein distance. SA-GDA [22] proposes a spectral augmentation module to enhance the node representation learning, which combines the target domain spectral information within the source domain. DMGNN [20] employes a GNN encoder with dual feature extractors to separate ego-embedding learning from neighbor-embedding learning and then a label propagation node classifier is employed to refine label prediction. DGDA [23] addresses graph domain adaptation in a generative view, which disentangles the generation process into the semantic latent variables, the domain latent variables, and the random latent variables.
@@
Despite the progress made by the aforementioned UGDA models, such solutions still struggle to address the complex domain shift present in real-world graph data. The complex distribution shift between graphs usually combine node attribute shift, graph structure shift and label shift. Thus, to deal with the distinct effects of distribution shifts caused by graph structures, StruRW [16] investigates different types of distribution shifts of graph-structured data and reweights edges in the source graph to reduces the conditional shift of neighborhoods. Based on this work, PairAlign [25] not only uses edge weights to recalibrate the influence among neighboring nodes to handle conditional structure shift but also adjusts the classification loss with label weights to handle label shift. Except for reweighting, JHGDA [17] designs a hierarchical pooling model to extract meaningful and adaptive hierarchical structures and jointly minimizes marginal and class conditional distribution shifts on each hierarchical level. KBL [18] redefines the aggregate mechanism as learning a knowledge-enhanced posterior distribution for target domains, which learns the scope of knowledge transfer by connecting knowledgeable samples between domains.
@@
Instead of simply using GNN as a node embedding module, some methods devote effort to empirically and theoretically studying the role of GNN within domain adaptation [14, 15, 24]. This allows the UGDA method to be precisely customized based on the property of GNN, thereby enhancing their performance in graph domain adaptation. GRADE [14] introduces graph subtree discrepancy as a metric to measure the graph distribution shift by connecting GNNs with WL subtree kernel [44]. SpecReg [15] finds the OT-based bound for graph is closely coupled with the Lipschitz constant of GNN and proposes spectral regularization to modulate the Lipschitz constant to restrict the target risk bound. A2GNN [24] further investigates the GNN’s underlying generalization capability behind its architecture and finds propagation operation plays a pivotal role. Based on this observation, A2GNN proposes a simple yet effective GNN which stacks more propagation layers on target branch. The timeline of UGDA algorithms covered in GDABench is shown in Figure 1 and more detailed descriptions are provided in Appendix C.
@@ 3 Datasets
We have carefully selected five widely used public datasets that showcase a wide spectrum of distribution shifts across graphs. These include Airport which consists of three domains Brazil (B), Europe (E) and USA (U); Blog that includes two domains Blog1 (B1) and Blog2 (B2); ArnetMiner which encompasses three domains DBLPv7 (D), Citationv1 (C) and ACMv9 (A); Twitch that includes six domains Germany (DE), England (EN), Spain (ES), France (FR), Porutgal (PT) and Russia (RU); MAG that includes six domains like CN, US, JP, FR, RU, and DE. The selection criteria for these datasets are primarily based on three factors: the complexity of the distribution shift, the scale of the dataset, and the potential for downstream applications. From these datasets, we have compiled a comprehensive collection comprising 74 distinct source-target adaptation pairs. Detailed attributes of each dataset are documented in Table 1 and the statistical methods for quantifying the types of domain shifts exhibited in the dataset is shown in Appendix B.
  • Wide range of distribution shift. Potential distribution shifts between source and target graphs can largely fall into three categories: feature shift, structure shift and label shift [31, 45, 25]. Our GDABench datasets capture a wide range of three distribution shifts across domains in varying degrees; The details are illustrated in Table 1 and more statistics are given in Appendix B. Specifically, the domains within Airport are dominated by structure shift, while domains in Blog, ArnetMiner, Twitch and MAG are affected by all kinds of shifts with different degrees.
  • Different scales with variant spans. We divide the size of the dataset by the average number of domain nodes. The small size (S) covers nodes below 5 thousand. The medium size (M) cover nodes below 10 thousand. The large size (L) covers nodes from 10 thousand to hundred thousand. Smaller datasets exhibit smaller differences in domain sizes, and vice versa. In the Blog and Airport, the size difference between the largest and smallest domains is less than 1000 nodes. In the case of ArnetMiner and Twitch, this difference ranges between 1,000 and 10,000. For the MAG, this difference ranges between 10,000 to 100,000. This allows us to understand the impacts of varying domain size on the efficacy of adaptation task.
  • Various downstream applications. The GDABench datasets encompass multiple applications, including citation relationships (ArnetMiner and MAG), social media (Blog and Twitch) and routines connections (Airport). Specifically, in ArnetMiner and MAG, nodes represent academic papers and edges indicate citation relationship. ArnetMiner groups domains by publisher, while MAG by country. Blog and Twitch capture friendship within blog and gamer networks, respectively. Airport delineates routines connections, where airports serve as nodes connected by flight routes.
@@ 4 Compared Models
Specialized UGDA Methods. This group includes specifically designed algorithms for graph domain adaptation task. We compare 16 models including (1) nine methods incorporating classicial DA methods with deep node embedding: DANE [9], ACDNE [10], UDAGCN [11], ASN [43], AdaGCN [12], CWGCN [21], SAGDA [22], DMGNN [20] and DGDA [23]; (2) four methods tailored for graph structure shift: StruRW [16], JHGDA [17], KBL [18] and PairAlign [25]; and (3) four methods based on domain adaptive message passing: GRADE [14], SpecReg [15], and A2GNN [24].
SimGDA: Vanilla DA with GNN Variants. To understand the inherent transferability of GNN, we delve into its aggregation process and decouple it from two perspectives: how to aggregate and what to aggregate. For how to aggregate, we consider five aggregator including sum aggregator [46], mean aggregator [47], aggregate with weighted neighbours (GAT) [48] and aggregate with discriminative neighbours (GIN) [49]. For what to aggregate, we consider three aspects in terms of the hop-count of neighbours: (1) GNN without neighbours, where graph structure is not considered (degenerating to MLP); (2) GNN with one-hop neighours; and (3) GNN with muti-hop neighours. To avoid the over-smooth problem, we also add residual connections to enhance its modeling power [50, 46]. For alignment, we consider two widely used models for domain-invariant feature learning from computer vision: domain distance metric MMD [34] and adversarial learning DANN [36]. MMD proposes to match the distribution in the latent space through maximum mean discrepancy [40]. DANN introduces an adversarial objective to distinguish source and target samples in the latent space. We use one-layer GCN [46] as a control and create six GNN variants by altering only one module each time. Then, we get 14 models by combining these variants with two vanilla DA methods, abbreviated these 14 models as SimGDA.
@@
SimGDA+: SimGDA with Unsupervised Techniques. To further unlock the power of GNN for graph domain adaptation, we enhance SimGDA with unsupervised graph learning techniques on unlabeled target graph, which allows the model to learn meaningful representations without relying on domain-specific labels. We implement three unsupervised techniques in an end-to-end manner: (1) Information Maximization (IM) [51, 52, 19]. Ideally, a perfect prediction for target domain should be individually certain and globally diverse. To achieve this goal, we minimize the entropy for each individual sample, and maximize the entropy for each class. (2) Graph AutoEncoder (AE) [53, 54, 55]. Graph autoencoders encode nodes into a latent vector space and reconstruct the graph data from the encoded latent space. After obtaining target graph node representations, we employ a distinct decoder to reconstruct target graph structure. (3) Graph Contrastive Learning (CL) [56, 57, 58, 59]: Graph contrastive learning methods maximize mutual information between augmented instances of the same object (e.g., node).
As changing the graph structure may affect the estimation of structure shift, we take random attribute masking to create augmented instances for each target node. Following previous works [57, 59], we utilize a one-layer perceptron (MLP) as projection head to map augmented representations to the shared latent space and then calculate contrastive loss based on normalized temperature-scaled cross entropy loss (NT-Xent) [60]. In total, we get 42 combined models, which integrates 14 SimGDA variants with 3 unsupervised techniques. We collectively refer to these combined models as SimGDA+, and the combination process is shown in Figure 2.

### v2

@@ Abstract
Unsupervised Graph Domain Adaptation (UGDA) involves the transfer of knowledge from a label-rich source graph to an unlabeled target graph under domain discrepancies. Despite the proliferation of methods designed for this emerging task, the lack of standard experimental settings and fair performance comparisons makes it challenging to understand which and when models perform well across different scenarios. To fill this gap, we present the first comprehensive benchmark for unsupervised graph domain adaptation named GDABench, which encompasses 16 algorithms across diverse adaptation tasks. Through extensive experiments, we observe that the performance of current UGDA models varies significantly across different datasets and adaptation scenarios. Specifically, we recognize that when the source and target graphs face significant distribution shifts, it is imperative to formulate strategies to effectively address and mitigate graph structural shifts. We also find that with appropriate neighbourhood aggregation mechanisms, simple GNN variants can even surpass state-of-the-art UGDA baselines. To facilitate reproducibility, we have developed an easy-to-use library PyGDA for training and evaluating existing UGDA methods, providing a standardized platform in this community. Our source codes and datasets can be found at https://github.com/pygda-team/pygda.
@@ 1 Introduction
The last decade has witnessed significant advancements in Graph Neural Networks (GNNs) with their successful applications spanning various fields [1, 2], including social network analysis [3, 4], protein interaction prediction [5], and traffic flow forecasting [6], etc. However, the presence of distribution shifts [7] and label scarcity in real-world graph data impedes the ability of existing GNN models to adapt to new domains [8]. To addresses this challenge, Unsupervised Graph Domain Adaptation (UGDA) has become an important solution for transferring knowledge from a labeled source graph to an unlabeled target graph. This powerful paradigm has been widely studied, unlocking broader application for graph neural networks.
@@
To fill these gaps, we revisit existing UGDA algorithms and conduct a comprehensive benchmark named GDABench. Specifically, GDABench includes 16 state-of-the-art UGDA models and diverse real-world graph datasets covering node attributes, graph structures, and label proportion shifts. Additionally, we also explore the limits of GNN transferability by combining 7 GNN variants with 2 domain alignment and 3 unsupervised graph learning techniques. Our work is the first to provide a rigorous empirical analysis of how various aggregation mechanisms influence alignments in domain adaptation task. Through comprehensive experiments, we observe that: (1) the performance of current UGDA models varies greatly across different datasets and adaptation scenarios; (2) it is crucial to develop tailored strategies to address graph structural shifts, especially when the distribution discrepancies are significant; (3) the GNN’s transferability in UGDA heavily relies on two factors: aggregation scope and aggregation architecture, which are influenced by the severity of label shift and the level of graph heterophily, etc; (4) the inherent adaptability of GNNs is largely underestimated by existing methods, which motivates the exploration of a simple yet effective model that fully leverages the core property of GNN. More insights can be found in Section 5.
@@
  • We introduce GDABench, the first comprehensive benchmark for unsupervised graph domain adaptation. It includes 16 recent state-of-the-art methods across various real-world datasets with diverse range of adaptation tasks.
@@ 2.1 Problem Definition
Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $n$ nodes and $m$ edges. The node feature matrix, denoted by $\mathbf{X}=\{{x_{v}|v\in}\mathcal{V}\}\in\mathbb{R}^{n\times d}$, contains attribute vector for each node, where $d$ represents the dimensionality of the attributes. We denote adjacency matrix as $\mathbf{A}\in\mathbb{R}^{n\times n}$, where $\mathbf{A}_{i,j}=1$ indicates the presence of an edge $e_{i,j}\in\mathcal{E}$ connecting node $v_{i}$ and $v_{j}$, and $\mathbf{A}_{i,j}=0$ otherwise. $\mathbf{Y}\in\mathbb{R}^{N\times C}$ represents the node label matrix, where $C$ is the category of node labels. We focus on Unsupervised Graph Domain Adaptation (UGDA) on classification tasks. It is a non-trivial task due to the complex domain shift between source and target graphs. Formally, given a labeled source graph $\mathcal{G}_{S}=(\mathcal{V}_{S},\mathcal{E}_{S},\mathcal{Y}_{S})$ and an unlabeled target graph $\mathcal{G}_{T}=(\mathcal{V}_{T},\mathcal{E}_{T})$ with the data shift that $\mathbb{P}_{S}(\mathcal{G})\neq\mathbb{P}_{T}(\mathcal{G})$. The goal is to train a graph neural network model $h:\mathcal{G}\rightarrow\mathcal{Y}$ that utilizes the labeled source graph $\mathcal{G}^{s}$ and the unlabeled target graph $\mathcal{G}^{t}$ in order to make accurate label predictions for the target graph $\mathcal{G}^{t}$.
@@ 2.2 Related Work
Classical Methods. There are two classical categories to reduce the domain discrepancy [33, 34]: minimizing pre-defined probability discrepancy metrics [35, 36] and adversarial learning techniques [37, 38, 39, 40]. For the pre-defined probability discrepancy metric minimization models, node representations are initially derived from the encoder, after which domain-invariant representations are acquired by minimizing probability discrepancy distances, such as MMD [41], CMD [36], etc. Rather than directly minimizing domain discrepancies, some approaches integrate the encoder with a domain classifier that predicts the source domain of each representation. For example, DANN [37] facilitates domain invariant learning (DIRL) to distinguish between source and target samples in the latent space. Building upon this framework, WDGRL [42] replaces the domain classifier with a network that learns an approximate Wasserstein distance. These classical methods are designed for CV and NLP tasks, where samples are independently and identically distributed [43]. However, marginal alignment of node representations in non-graph DA research is insufficient for graph-structured data, of which the distribution shift becomes more complex due to the interconnection among different nodes.
@@
Specialized Methods for Graph Data. To tackle the unique challenges of knowledge transfer in graphs, several approaches have been proposed [9, 11, 44, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. The key idea behind existing work on node-level UGDA is to leverage node representations as intermediaries while minimizing domain shift through adversarial learning. This kind of methods incorporate classical DA methods with deep node embedding by designing a special encoder to learn transferable features. DANE [10] uses shared weight GCNs to get node representations and then handles distribution shift via least square generative adversarial network. ACDNE [11] employs two feature extractors to simultaneously maintain both attributed affinity and topological proximity through a deep network embedding module. Additionally, it integrates a domain classifier to enhance the label-discriminative nature of the node representations. UDAGCN [12] introduces a dual graph convolutional network enhanced by an attention mechanism, allowing it to leverage both local and global consistency for improved graph representation learning. ASN [44] distinguishes between domain-private and domain-shared information while integrating both local and global consistency to effectively capture network topology information. AdaGCN [13] utilizes GCNs to merge network topology with adversarial domain adaptation, effectively enhancing graph convolution processes. CWGCN [22] introduces a two-step correntropy-induced Wasserstein GCN. This method first eliminates noisy nodes from the source graph and subsequently learns the target GCN by extending the Wasserstein distance. SA-GDA [23] introduces a spectral augmentation module aimed at improving node representation learning by integrating spectral information from the target domain into the source domain. DMGNN [21] utilizes a GNN encoder equipped with dual feature extractors to distinguish between ego-embedding learning and neighbor-embedding learning. Following this, a label propagation node classifier is applied to enhance the accuracy of label predictions. DGDA [24] approaches graph domain adaptation from a generative perspective, breaking down the generation process into three components: semantic latent variables, domain latent variables, and random latent variables.
@@
Despite the progress made by the aforementioned UGDA models, such solutions still struggle to address the complex domain shift present in real-world graph data. The complex distribution shift between graphs usually combine node attribute shift, graph structure shift and label shift. Thus, to deal with the distinct effects of distribution shifts caused by graph structures, StruRW [17] investigates different types of distribution shifts of graph-structured data and reweights edges in the source graph to reduce the conditional shift of neighborhoods. Based on this work, PairAlign [26] addresses conditional structure shifts by recalibrating the influence of neighboring nodes using edge weights, while also modifying the classification loss through label weights to tackle label shifts. Except for reweighting, JHGDA [18] develops a hierarchical pooling model that extracts meaningful and adaptive structures at various levels. This model simultaneously minimizes both marginal and class conditional distribution shifts across each hierarchical layer. KBL [19] redefines the aggregation mechanism as a process of learning a knowledge-enhanced posterior distribution for target domains, facilitating knowledge transfer by linking informative samples across domains.
@@
Instead of simply using GNN as a node embedding module, some methods devote effort to empirically and theoretically studying the role of GNN within domain adaptation [15, 16, 25]. This allows the UGDA method to be precisely customized based on the property of GNN, thereby enhancing their performance in graph domain adaptation. GRADE [15] introduces graph subtree discrepancy as a metric to measure the graph distribution shift by connecting GNNs with WL subtree kernel [45]. SpecReg [16] finds that the OT-based bound for graph is closely coupled with the Lipschitz constant of GNN and proposes spectral regularization to modulate the Lipschitz constant to restrict the target risk bound. A2GNN [25] further investigates the GNN’s underlying generalization capability behind its architecture and finds propagation operation plays a pivotal role in the adaptation procedure. Based on this observation, A2GNN proposes a simple yet effective GNN framework, which stacks more propagation layers on target branch. The timeline of UGDA algorithms covered in GDABench is shown in Figure 1 and more detailed descriptions are provided in Appendix C.
@@ 3 Datasets
We have carefully selected 5 widely used public datasets that showcase a wide spectrum of distribution shifts across graphs for the node classification task. These include Airport which consists of three domains: Brazil (B), Europe (E) and USA (U); Blog that includes two domains: Blog1 (B1) and Blog2 (B2); ArnetMiner which encompasses three domains: DBLPv7 (D), Citationv1 (C) and ACMv9 (A); Twitch that includes six domains: Germany (DE), England (EN), Spain (ES), France (FR), Porutgal (PT) and Russia (RU); MAG that includes six domains like CN, US, JP, FR, RU, and DE. The selection criteria for these datasets are primarily based on three factors: the complexity of the distribution shift, the scale of the dataset, and the potential for downstream applications. From these datasets, we have included a comprehensive collection, comprising 74 distinct source-target adaptation pairs. Detailed information about each dataset is provided in Table 1, while the statistical methods used to quantify the types of domain shifts exhibited in the dataset are presented in Appendix B. The chosen datasets possess the following characteristics:
  • Wide range of distribution shift. The graph distribution shifts between source and target domains can largely fall into three categories: feature shift, structure shift and label shift [32, 46, 26]. Our GDABench datasets encompass a diverse range of three distribution shifts across domains in varying degrees. The details are illustrated in Table 1 and more statistics are given in Appendix B. Specifically, the domains within Airport are dominated by structure shift, while domains in Blog, ArnetMiner, Twitch and MAG are affected by all kinds of shifts with different degrees.
  • Different scales with variant spans. We categorize the size of the dataset by the average number of domain nodes. The small size (S) covers nodes below 5 thousand. The medium size (M) cover nodes below 10 thousand. The large size (L) covers nodes from 10 thousand to hundred thousand. Smaller datasets exhibit smaller differences in domain sizes, and vice versa. In the Blog and Airport, the size difference between the largest and smallest domains is less than 1000 nodes. In the case of ArnetMiner and Twitch, this difference ranges between 1,000 and 10,000. For the MAG, this difference ranges between 10,000 to 100,000. This allows us to understand the impacts of varying domain size on the efficacy of adaptation task.
  • Various downstream application scenarios. The GDABench datasets encompass multiple application scenarios, including citation relationships (ArnetMiner and MAG), social media interactions (Blog and Twitch) and routine connections (Airport). Specifically, in ArnetMiner and MAG, nodes represent academic papers and edges indicate citation relationship. ArnetMiner groups domains by publisher, while MAG by country. Blog and Twitch capture friendship within blog and gamer networks, respectively. Airport delineates routines connections, where airports serve as nodes connected by flight routes.
@@ 4 Compared Models
Specialized UGDA Methods. This group includes specifically designed algorithms for graph domain adaptation task. We compare 16 models including (1) nine methods incorporating classicial DA methods with deep node embedding: DANE [10], ACDNE [11], UDAGCN [12], ASN [44], AdaGCN [13], CWGCN [22], SAGDA [23], DMGNN [21] and DGDA [24]; (2) four methods tailored for graph structure shift: StruRW [17], JHGDA [18], KBL [19] and PairAlign [26]; and (3) three methods based on domain adaptive message passing: GRADE [15], SpecReg [16], and A2GNN [25].
SimGDA: Vanilla DA with GNN Variants. To understand the inherent transferability of GNN, we delve into its aggregation process by decoupling it into two key perspectives: how to aggregate and what to aggregate. For how to aggregate, we consider five types of aggregators, including sum aggregator [47], mean aggregator [48], aggregate with weighted neighbours (GAT) [49] and aggregate with discriminative neighbours (GIN) [50]. For what to aggregate, we consider three aspects in terms of the hop-count of neighbours: (1) GNN without neighbours, where graph structure is not considered (degenerating to MLP); (2) GNN with one-hop neighours; and (3) GNN with multi-hop neighours. To avoid the over-smooth problem, we also add residual connections to enhance its modeling power [51, 47]. For alignment, we consider two widely used models for domain-invariant feature learning from computer vision: domain distance metric MMD [35] and adversarial learning DANN [37]. Among them, MMD proposes to match the distribution in the latent space through maximum mean discrepancy [41], while DANN introduces an adversarial objective to distinguish source and target samples in the latent space. We use one-layer GCN [47] as a control and create six GNN variants by altering only one module each time. Then, we get 14 models by combining these variants with two vanilla DA methods, abbreviated these 14 models as SimGDA.
@@
SimGDA+: SimGDA with Unsupervised Techniques. To further unlock the power of GNN for graph domain adaptation, we enhance SimGDA with different unsupervised graph learning techniques on unlabeled target graph, which allows the model to learn meaningful representations without relying on domain-specific labels. We implement three unsupervised techniques in an end-to-end manner: (1) Information Maximization (IM) [52, 53, 20]. Ideally, an accurate prediction for the target domain should exhibit individual certainty while maintaining global diversity. To accomplish this, we minimize the entropy for each individual sample while maximizing the entropy across classes. (2) Graph AutoEncoder (AE) [54, 55, 56]. Graph autoencoders encode nodes into a latent vector space and reconstruct the graph data from the encoded latent space. After obtaining target graph node representations, we employ a distinct decoder to reconstruct target graph structure. (3) Graph Contrastive Learning (CL) [57, 58, 59, 60]: Graph contrastive learning methods maximize mutual information between augmented instances of the same object (e.g., node).
As changing the graph structure may affect the estimation of structure shift, we take random attribute masking to create augmented instances for each target node. Following previous works [58, 60], we utilize a one-layer perceptron (MLP) as projection head to map augmented representations to the shared latent space and then calculate contrastive loss based on normalized temperature-scaled cross entropy loss (NT-Xent) [61]. As a result, we get 42 combined models, which integrates 14 SimGDA variants with 3 unsupervised techniques. We collectively refer to these combined models as SimGDA+, and the combination process is shown in Figure 2.